OpenStudy (vanz):

Whats a real world example for radicals? Like in what situation would I use it?? (liberal arts)

OpenStudy (nottim):

What radicals? Math?

OpenStudy (vanz):

yes.

OpenStudy (nottim):

Why the liberal arts?

OpenStudy (vanz):

bc thats a math class, its jsut in case anyone ask for what class i need it for.

OpenStudy (nottim):

Makes sense.

OpenStudy (nottim):

Also, if Riddler ever asked a math question involving math, you would be Batman to the location of the hostages.

OpenStudy (nottim):

Beat, not be. You can't be Batman.

hero (hero):

Popular Uses For Radicals Radius Of A Cylinder: \[V = \sqrt{\frac{V}{\pi h}}\] Free-Falling Object: \[t = \sqrt{\frac{2s}{g}} \] Current: \[I = \sqrt{\frac{P}{R}}\] Geometric Mean \[x = \sqrt{ab} \] Period Of Pendulum: \[T = 2\pi \sqrt{\frac{l}{9.8}}\]

OpenStudy (nottim):

And you would use physics in real life. Who wouldn't like to calculate events before, during or after events?

hero (hero):

I goofed up on the first one. The V on the left should be r

OpenStudy (vanz):

Thank you so much! also when would I use factoring in real life??

OpenStudy (nottim):

What's factoring.

OpenStudy (nottim):

Hello?

OpenStudy (nottim):

You good?

OpenStudy (vanz):

oh sorry, hold on let me check :D

OpenStudy (vanz):

I just need help with one more!

OpenStudy (vanz):

Whats a real world example for quadratics?

OpenStudy (nottim):

WHOA. YOU loading this up a notch! Lemme find one.

OpenStudy (nottim):

Both are long reads, sorry. Do a skim thru tho.

OpenStudy (vanz):

will do!! thanks for all your help!!

OpenStudy (nottim):

Your welcome. My work here is done. BATMAN AWAY